Needing to write and solve equation that will enable me to determine the length and width of the rectangle. L=6cm longer than its width
P=32cm

Answer by rvfer(117)

2(x+6)+2x=32
2x+12+2x=32
4x=32-12
4x=20
x=20/4
x=5
x+6=11